# Optimal Virtual Model Control for Robotics: Design and Tuning of Passivity-Based Controllers

### v2

@@ VII Conclusions
Our approach has two fundamental limitations. One is intrinsic to the non-convexity of the optimization problem (13), leading to local minima and lack of convergence guarantees. However, for the parameterization and cost functions of our examples, we have shown that the optimization converges, recovering classical linear results (Section V-A), and handling the complexity of the Franka-Emika manipulator (Section V-B). The other limitation is the dependence of performance optimization on the accuracy of the robot model, which is challenging, particularly with phenomena such as non-linear friction dependent on load / temperature [52] and transmission non-linearities [53, 54].
@@
In the limit of highly parameterized and highly numerous components, our approach looks similar to physics-informed neural networks, as seen in [34] where the controller is a passive highly parameterized neural network. The trade-off between the flexibility of the controller for performance optimization and the physical intuition provided by virtual mechanisms can be explored by placing highly parameterized components at specific points on the virtual mechanism. Furthermore, the adoption of highly parameterized structural parts can further improve performance while retaining the physical intuition on the behavior of the controller / virtual mechanism. Future research will focus on these important aspects. We will also look at data-driven tuning to reduce the gap between simulation and experimental performances.
Acknowledgments

### v3

@@ VIII Conclusions
Our approach has two fundamental limitations. One is intrinsic to the non-convexity of the optimization problem (19), leading to local minima and lack of convergence guarantees. However, for the parameterization and cost functions of our examples, we have shown that the optimization converges, recovering classical linear results (Section VI-A), and handling the complexity of the Franka-Emika manipulator (Section VI-B). The other limitation is the dependence of performance optimization on the accuracy of the robot model, which is challenging, particularly with phenomena such as non-linear friction dependent on load / temperature [Bittencourt2012a] and transmission non-linearities [Kennedy2003, Ma2018].
@@
Optimal virtual model control could be used to help design structures linking leader/follower manipulators, with an adaptation of the cost function to optimise for teleoperative transparency. Alternatively, the automatic differentiation approach presented here could be used to optimise the performance in model-mediated teleoperation [Xu2016].
In the limit of highly parameterized and highly numerous components, our approach looks similar to physics-informed neural networks, as seen in [Massaroli2022] where the controller is a passive highly parameterized neural network. The trade-off between the flexibility of the controller for performance optimization and the physical intuition provided by virtual mechanisms can be explored by placing highly parameterized components at specific points on the virtual mechanism. Furthermore, the adoption of highly parameterized structural parts can further improve performance while retaining the physical intuition on the behavior of the controller / virtual mechanism. Future research will focus on these important aspects. We will also look at data-driven tuning to reduce the gap between simulation and experimental performances.
[Uncaptioned image] Daniel Larby received the MEng degree in control and information engineering in 2020 from the University of Cambridge, UK. He completed his Ph.D. degree in 2025, focussing on robotic control with applications in robotic surgery with the control lab, also in the University of Cambridge, UK. He now work for Swan EndoSurgical, part of a team working to develop robotic endoluminal surgery solutions. His research interests include robotics, virtual-mechanism control, passivity based control, impedance control, robotic surgery, and algorithmic differentiation.
[Uncaptioned image] Fulvio Forni Fulvio Forni is a Professor of Control Engineering at the University of Cambridge, where he has been a faculty member since October 2015. He earned his PhD from the University of Rome ‘Tor Vergata’ in 2010. Following his doctorate, he conducted postdoctoral research at the University of Liège in Belgium. Forni’s research interests encompass feedback control and robotics. He received the prestigious IEEE CSS George S. Axelby Outstanding Paper Award in 2020. He is a Director of Studies of Newnham College at Cambridge and serves as a co-investigator for the EPSRC Centre for Doctoral Training in Agrifood Robotics ‘Agriforwards’.